# Efficient AI in Practice: Training and Deployment of Efficient LLMs for Industry Applications

### v1

@@ A.1. Comparison of Quantization Methods
We present a comparison of different quantization methods. We use the Meta Llama 3.1 8B Instruct model, and quantize the model using 1024 samples from the open source C4 (Raffel et al., 2020) dataset as the calibration set. We report the zero-shot accuracy of the model on three open-source tasks PIQA (Bisk et al., 2020) and ARC easy/challenge (Clark et al., 2018). We compare W8A8 quantization with SmoothQuant (Xiao et al., 2023), FP8 quantization on H100 GPUs, and W4A16 quantization with GPTQ (Frantar et al., 2022) and QuantEase (Behdin et al., 2023). The results are shown in Table 8. We see that 8-bit quantization generally has a small loss of accuracy. On the other hand, GPTQ with W4A16 shows some model quality degradation. However, using QuantEase for better optimization helps to reduce the model quality gap. In our internal experiments, we have observed similar trends when comparing different methods.
@@
Table 8. Comparison of different quantization schemes with the Llama 3.1 8B Instruct model.
A.2. Extra tables for Section 5
@@
Table 9. Results for $m=1$, $k=1$ for $16k$ context length using 4 GPUs (tp=4).
@@
Table 10. Results for $m=1$, $k=1$ for $32k$ context length using 4 GPUs (tp = 4).
@@
Table 11. Results for $m=1$, and $k=4$ for $32k$ context length using 4 GPUs (tp=4).

### v2

@@ B.1 Comparison of Quantization Methods
We present a comparison of different quantization methods. We use the Meta Llama 3.1 8B Instruct model, and quantize the model using 1024 samples from the open source C4 Raffel et al. (2020) dataset as the calibration set. We report the zero-shot accuracy of the model on three open-source tasks PIQA Bisk et al. (2020) and ARC easy/challenge Clark et al. (2018). We compare W8A8 quantization with SmoothQuant Xiao et al. (2023), FP8 quantization on H100 GPUs, and W4A16 quantization with GPTQ Frantar et al. (2022) and QuantEase Behdin et al. (2023). The results are shown in Table 7. We see that 8-bit quantization generally has a small loss of accuracy. On the other hand, GPTQ with W4A16 shows some model quality degradation. However, using QuantEase for better optimization helps to reduce the model quality gap. In our internal experiments, we have observed similar trends when comparing different methods.
@@
Table 7: Comparison of different quantization schemes with the Llama 3.1 8B Instruct model.
Full PrecisionC4 1024C4 4096Task-Specific 350Task-Specific 7001-10.8-0.80.6-0.60.4-0.40.2-0.2001.06-1.060.64-0.640.12-0.120.1-0.1AUC Delta (%)
Figure 5: Comparison of one-shot pruning methods. The bars indicate the drop (in percentage points) relative to the full precision baseline. The pruned model is a 6.4B model (20% MLP pruning).
B.2 Effect of calibration data for Pruning
Figure 5 captures the effect of the calibration dataset $\boldsymbol{X}$ (discussed in Section A and Equation (4)) on the accuracy of the pruned versions of the 8B student model (results for the 3B model are similar and are hence omitted for clarity). The Full Precision bar illustrates the baseline accuracy of the non-pruned model. We consider two different datasets for calibration - C4 Raffel et al. (2020), an open source dataset, and an in-domain dataset. When we prune using a randomly sampled portion of the C4 dataset (1,024 or 4,096 examples), accuracy drops, although more samples mitigate this drop to an extent. These results indicate that increasing the number of calibration examples from 1,024 to 4,096 can partially recover lost accuracy due to pruning. However, leveraging fewer but more domain-specific samples (350 or 700 examples from the target task) yields better accuracy values, which closely match the full precision baseline. This highlights the importance of using task-relevant data for calibration, even if it involves fewer examples, as it can produce more accurately pruned models than generic calibration sets.
B.3 Extra tables for Section 6
@@
Table 8: Results for $m=1$, $k=1$ for $16k$ context length using 4 GPUs (tp=4).
@@
Table 9: Results for $m=1$, $k=1$ for $32k$ context length using 4 GPUs (tp = 4).
@@
Table 10: Results for $m=1$, and $k=4$ for $32k$ context length using 4 GPUs (tp=4).
Appendix C Implementation and Efficiency Checklist
We provide an open source implementation of all methods discussed in this paper, in addition to examples to recreate our distillation and pruning pipelines. Our python package can be found at https://github.com/linkedin/FMCHISEL.
Additionally, we summarize the lessons from our experiments by providing a practical checklist to create efficient SLMs through distillation and pruning for a specific task. We assume one has access to a family of pre-trained LLMs such as Llama.
  1. 1.
    Create a teacher model, for example, by fine-tuning a large pre-trained on the desired task(s).
  2. 2.
    Perform distillation using the teacher model developed above, and a variety of student models.
  3. 3.
    Find the smallest student model that meets the desired quality bar.
  4. 4.
    Perform profiling on the selected student model to identify inference bottlenecks.
  5. 5.
    Prune and distill the selected student model, targeting bottlenecks identified above, until the pruned model’s quality approaches the quality bar. This is the final model that will be deployed.
  6. 6.
    Perform extensive benchmarking under various workloads and scenarios (such as hardware failure, bursty traffic, etc.) to obtain a conservative estimation of the number of GPUs required.
Appendix D Ethical Concerns
This paper presents a methodological study of model compression techniques such as model pruning and knowledge distillation. These techniques aim to ensure the compressed models produce similar predictions compared to the original ones and hence, these methods do not pose any inherent ethical concerns. The analysis of larger (teacher) models is out of the scope of this paper.